OpenStudy (anonymous):

Provide an example of an arithmetic series which totals zero. Using complete sentences, explain how you created the example.

OpenStudy (anonymous):

sum of AP = n/2{2a+(n-1)d} =0. where a= first term , d=common diff , n =numbers of term. now it gives (n-1)d = -2a. infinite numbers of values of n,d , and a satisfy the equation so infinite no of series can be found... for examle put n=10 ,d=2 then a=-9.. then the series.....-9,-7,-5,-3 ............ in similar manner you can get infinite no of series by putting different values of a, n n and d which satisfy the equation..